OpenStudy (anonymous):

An investment is worth $3518 in 1995. By 2000 it has grown to $5553. Let y be the value of the investment in the year x, where x = 0 represents 1995. Write a linear equation that models the value of the investment in the year x.

OpenStudy (paxpolaris):

linear equation: \(\large y=mx+b\) at x= 0 y=3518... \(\therefore b=3518\) \[\large m={5553-3518\over5-0}\]

OpenStudy (anonymous):

here are the choices A. y = -407x + 7588 B. y = x + 3518 C. y = -407x + 3518 D. y = 407x + 3518

OpenStudy (paxpolaris):

\[\large m={5553-3518\over5-0}={2035 \over5}=\huge 407\] b=3518 y=mx+b so, y= 407x+3518

OpenStudy (paxpolaris):

D.

OpenStudy (anonymous):

thnxx. well explained!